Recently, Midrand Hospitals Inc. filed for bankruptcy. The firm was reorganized as American Hospitals Inc., and the court permitted a new indenture on an outstanding bond issue to be put into effect. The issue has 10 years to maturity and a coupon rate of 10 percent, paid annually. The new agreement allows the firm to pay no interest for 5 years. Then, interest payments will be resumed for the next 5 years. Finally, at maturity (Year 10), the principal plus the interest that was not paid during the first 5 years will be paid. However, no interest will be paid on the deferred interest. If the required annual return is 20 percent, what should the bonds sell for in the market today?
1
Expert's answer
2020-09-02T18:25:07-0400

year 1-5: $0 per annum


year6-10: $100 per annum coupon payments


Year 10:Par value+deferred interest for the 1st 5 years


=1,000+(5"\times"100) =1,500


The required rate of return is to 0.2% per annum


The PV of the cash-flows will be ;

PV of the coupon payments + PV of the par value plus deffred interest


=100"\times"PV Annuity Factor for 5 periods at 20%"\times"PV Interest factor with i=20% and n =5$ plus 1,500×PV Interest factor with i=20% and n =10


Thus the bonds should sell for ;


= "100\times \frac{[1-(1+0.2)]}{0.2(1+0.2)^{-5} }+\frac{1,500}{(1+0.2)^{10}}"


= $362.4


Thus the bonds should sell for $362.4 in the market today.